Show that A(3,5),B(-3,-2), and C(3,-8) are the vertices of an isosceles triangle
Found 2 solutions by Alan3354, Boreal:
Answer by Alan3354(69443)

If the 3 side lengths are equal, it is.

Answer by Boreal(15235)

Show that the distance of two of the sides is the same
AB has sqrt (49+36)=sqrt (85)
BC has sqrt (36+36)=6 sqrt (2)
CA has sqrt (169)=13
These are unequal sides.